How many litres of water must be added to 1 litre of an aqueous solution of HCl with a pH of 1 to create an aqueous solution with pH of 2? (1) 0.9 L (2) 2.0 L (3) 9.0 L (4) 0.1 L

Grade:upto college level

3 Answers

Sunil Kumar FP
9 years ago
Initial pH = 1, i.e. [H+] = 0.1 mole/litre
New pH = 2, i.e. [H+] = 0.01 mole/litre
In case of dilution: M1V1= M2V2
0.1 ×1 =0.01 × V2
V2= 10 litre.
Volume of water added = 9 litre.
Hence (3) is correct.
thanks and regards
sunil kr
askIITian faculty
Ovais
5 years ago
HCl ph=1,v=1
H+=1/10
HCl ph=2,v=?
H+=1/100
M1v1=M2V2
1/10×1=1/100×V2
V2=10
Hence solved
If u can solve more easily plz answer
Vf=v2_V1
Vf=10_1=9l
Yash Chourasiya
3 years ago
Dear Student

Initial pH = 1, i.e. [H+] = 0.1 mole/litre
New pH = 2, i.e. [H+] = 0.01 mole/litre
In case of dilution: M1V1 = M2V2
0.1 ×1 =0.01 × V2
V2= 10 litre.
Volume of water added = 10 – 1 = 9 litre

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya